# Asymmetric Perturbation in Solving Bilinear Saddle-Point Optimization

### v1

@@ 4.1 Last-iterate convergence rate
In this section, we provide the convergence result of AsymP-GDA. Let $\|A\|$ denote the largest singular value of a given matrix $A$, and $D:=\max_{z,z^{\prime}\in\mathcal{X}\times\mathcal{Y}}\left\|z-z^{\prime}\right\|$ denote the diameter of $\mathcal{X}\times\mathcal{Y}$. Theorem 4.1 demonstrates that AsymP-GDA converges to the minimax/maximin strategies in the asymmetrically perturbed game at a rate of $\mathcal{O}(1/t)$:
@@
For an arbitrary perturbation strength $\mu>0$, if the learning rate satisfies $\eta<\min\left(\frac{\mu}{2(\mu^{2}+\|A\|^{2})},\frac{8\left(\mu+\|A\|\right)}{D\min\left(\mu,\frac{\beta}{\mu}\right)}\right)$, then $x^{t}$ and $y^{t}$ satisfy for any $t\geq 1$:
$\left\|x^{\mu}-x^{t}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t})-y^{t}\right\|^{2}\leq\frac{256D^{2}(\mu+\|A\|)^{2}}{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)t},$
where $\beta>0$ is a positive constant depending only on $\mathcal{Y}^{\mu}$.
Note that the statement of Theorem 4.1 holds for any fixed $\mu>0$. Thus, by combining Theorems 3.3 and 4.1, we can conclude that if $\mu$ is sufficiently small (which does not need to depend on the number of iterations $t$ or $T$), player $x$’s strategy $x^{t}$ updated by AsymP-GDA converges to a minimax strategy in the original game (1):
@@
Assume that the perturbation strength $\mu$ is set such that $\mu\in(0,\frac{\alpha}{\max_{x^{\ast}\in\mathcal{X}^{\ast}}\left\|x^{\ast}\right\|})$, and the learning rate is set such that $\eta<\min\left(\frac{\mu}{2(\mu^{2}+\|A\|^{2})},\frac{8\left(\mu+\|A\|\right)}{D\min\left(\mu,\frac{\beta}{\mu}\right)}\right)$. Then, AsymP-GDA ensures the convergence of $x^{t}$ to an equilibrium $x^{\ast}$ in the original game at a rate of $\mathcal{O}(1/t)$:
$\left\|x^{\ast}-x^{t}\right\|^{2}\leq\frac{256D^{2}(\mu+\|A\|)^{2}}{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)t}.$
Remark 4.3 (Overcoming the limitation of small fixed $\mu$).
One might still be concerned that the permissible perturbation strength $\mu$ in Corollary 4.2 must remain very small when $\alpha$ is close to 00. While it is true that $\alpha$ can, in theory, become arbitrarily small depending on the game’s payoff structure, such pathological cases are uncommon in practice, as demonstrated by our experiments. To mitigate this potential issue, a simple anchoring mechanism can be employed. This approach replaces the perturbation term $\frac{\mu}{2}\|x\|^{2}$ with $\frac{\mu}{2}\|x-\sigma\|^{2}$, where the anchor point $\sigma$ is periodically updated to the current strategy $x^{t}$ (Perolat et al., 2021; Abe et al., 2024). As discussed in AppendixC, this modification keeps the convergence guarantee and remains effective even when $\mu$ is moderately large.
@@
Firstly, leveraging the strong convexity of the perturbation payoff function, $\frac{\mu}{2}\left\|x\right\|^{2}$, we can show that the distance between the current strategy profile $z^{t}=(x^{t},y^{t})$ and any equilibrium $z^{\mu}=(x^{\mu},y^{\mu})$ monotonically decreases. Specifically, we have for any $t\geq 1$:
$\left\|z^{\mu}-z^{t+1}\right\|^{2}-\left\|z^{\mu}-z^{t}\right\|^{2}\leq-\eta\mu\left\|x^{\mu}-x^{t+1}\right\|^{2}-\frac{1}{2}\left\|z^{t+1}-z^{t}\right\|^{2}.$ (7)
@@
The primary technical challenge is deriving the term related to the distance between $y^{t+1}$ and the maximin strategies set $\mathcal{Y}^{\mu}$, i.e., $\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}$, which leads to the last-iterate convergence rate. To this end, we first derive the following lower bound on the path length $\left\|z^{t+1}-z^{t}\right\|$ by the distance $\|x^{\mu}-x^{t+1}\|^{2}$ and $\|A(y^{\mu}-y^{t+1})\|^{2}$ (as shown in Lemmas F.3 and F.4):
$\left\|z^{t+1}-z^{t}\right\|$ $\geq\Omega\left(\left\|x^{\mu}-x^{t+1}\right\|^{2}+\left\|A(y^{\mu}-y^{t+1})\right\|^{2}\right).$ (8)
One might think that the term of $\left\|A(y^{\mu}-y^{t+1})\right\|^{2}$ can be 00 even if $y^{t+1}\notin\mathcal{Y}^{\mu}$. However, by proving that $Ay^{\ast}$ attains a unique vector $b^{\ast}\in\mathbb{R}^{m}$ for all $y^{\ast}\in\mathcal{Y}^{\mu}$, and this vector cannot be achieved by any strategy $y\notin\mathcal{Y}^{\mu}$, we demonstrate that $\left\|A(y^{\mu}-y^{t+1})\right\|^{2}\neq 0$ as long as $y\notin\mathcal{Y}^{\mu}$. Using this fact, we demonstrate that the term of $\left\|A(y^{\mu}-y^{t+1})\right\|^{2}$ can be lower bounded by the distance between $y^{t+1}$ and the maximin strategy set:
$\left\|A(y^{\mu}-y^{t+1})\right\|^{2}\geq\Omega\left(\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}\right).$ (9)
By combining (8) and (9), we obtain the following lower bound on $\left\|z^{t+1}-z^{t}\right\|$ by the distance between the current strategy profile and the equilibrium set:
$\!\left\|z^{t+1}\!-\!z^{t}\right\|\!\geq\!\Omega\left(\!\left\|x^{\mu}\!-\!x^{t+1}\right\|^{2}\!+\!\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})\!-\!y^{t+1}\right\|^{2}\right)\!.$ (10)
@@
$\left\|\prod_{\mathcal{Z}^{\mu}}(z^{t+1})-z^{t+1}\right\|^{2}\leq\left\|\prod_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}-\Omega\left(\left\|\prod_{\mathcal{Z}^{\mu}}(z^{t+1})-z^{t+1}\right\|^{4}\right),$
where $\mathcal{Z}^{\mu}:=\{x^{\mu}\}\times\mathcal{Y}^{\mu}$. Finally, utilizing an auxiliary lemma on recursive formulas by (Wei et al., 2021), we obtain the following upper bound on the distance $\left\|\prod_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}$:
$\left\|\prod_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}\leq\mathcal{O}(1/t).$
@@
The main technical challenge in proving Theorem 4.1 arises from the asymmetric nature of the perturbation, which is applied only to player $x$. Unlike the symmetric case, where strong convexity in both players’ payoff functions directly yields contraction, the asymmetric setting requires a more subtle analysis since player $y$’s payoff function remains linear. Our proof establishes a last-iterate convergence rate by lower bounding the projection error of $y^{t}$ through the gap in payoff vectors $Ay^{t}-Ay^{\mu}$, as shown in (9), rather than relying on strong convexity. This contrasts with techniques for unconstrained linear-quadratic games (Zhang et al., 2022), where the dynamics reduce to a linear system; in our constrained setting, projections alter the dynamics fundamentally. See Appendix F for details.
5 Asymmetrically perturbed CFR+
We introduce the concept of asymmetric perturbation into learning algorithms for extensive-form games, which involve sequential decision-making and imperfect information, making them a more complex and realistic setting for strategic interactions. Here, we provide only empirical simulations, unlike in Section 4, as analyzing extensive-form games is significantly more difficult than analyzing normal-form games.
Extensive-form games are commonly solved using Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007) and its variants, such as CFR+ (Tammelin, 2014). In CFR+, the cumulative counterfactual regret $\boldsymbol{R}^{x}$ of player $x$ is updated to ensure it does not fall below zero:
$\boldsymbol{R}^{x}_{t+1}$ $=\left[\boldsymbol{R}^{x}_{t}+\boldsymbol{r}({x}_{t})\right]^{+},$
where $\boldsymbol{r}({x}_{t})$ represents the immediate regret incurred at iteration $t$. To investigate the impact of perturbations in extensive-form games, we propose Asymmetric Perturbed CFR+ (AsymP-CFR+), which incorporates an asymmetric perturbation, alongside Symmetric Perturbed CFR+ (SymP-CFR+) as a baseline, which applies a symmetric perturbation. In AsymP-CFR+, the cumulative counterfactual regret $\boldsymbol{R}^{x}$ for player $x$ is updated using the following rule:
$\boldsymbol{R}^{x}_{t+1}=\left[\boldsymbol{R}^{x}_{t}+\boldsymbol{r}({x}_{t})-\mu\boldsymbol{x}_{t}\right]^{+}.$
In contrast, SymP-CFR+ applies perturbation symmetrically to both players. The pseudocode for AsymP-CFR+ is provided in Algorithm 1 of Appendix A.4.
@@ 5 Asymmetrically perturbed CFR+
We compare the NashConv of the last-iterate strategy for AsymP-CFR+ against SymP-CFR+ and baseline algorithms, including CFR, CFR+, LCFR, and DCFR (both from Brown and Sandholm (2019)). Our experiments focus on six different extensive-form games: Kuhn Poker, Leduc Poker, Goofspiel (with four-card and five-card variants), and Liar’s Dice (with four-sided and six-sided dice), all of which are implemented using OpenSpiel (Lanctot et al., 2019). For both AsymP-CFR+ and SymP-CFR+, we set $\mu=0.01$. For DCFR, we use OpenSpiel’s default parameters.
@@
Figure 4 shows the NashConv values for each game. As indicated by these results, AsymP-CFR+ not only converges faster than any other method in all games but also directly reaches an equilibrium strategy, whereas SymP-CFR+ converges near the equilibrium. These results confirm that the asymmetric perturbation leads to convergence in extensive-form games.
Figure 4: Performance of AsymP-CFR+ in extensive-form games.

### v2

@@ 4.1 Last-Iterate Convergence Rate
In this section, we provide a last-iterate convergence result for AsymP-GDA. Let $\|A\|$ denote the largest singular value of the matrix $A$. As discussed in the previous section, when both players’ objectives are perturbed by strongly convex penalties, standard first-order methods converge to an approximate equilibrium at a linear rate (Cen et al., 2021, 2023; Pattathil et al., 2023; Sokota et al., 2023). Theorem 4.1 shows that, perhaps surprisingly, a linear convergence rate can still be achieved even when the perturbation is applied only to one player:
@@
For an arbitrary perturbation strength $\mu>0$, if the learning rate satisfies $\eta\leq\frac{\mu}{\mu^{2}+\|A\|^{2}}$, then $x^{t}$ and $y^{t}$ satisfy:
$\left\|\Pi_{\mathcal{Z^{\mu}}}(z^{t})-z^{t}\right\|^{2}\leq\left\|\Pi_{\mathcal{Z^{\mu}}}(z^{1})-z^{1}\right\|^{2}\cdot\exp(-\Theta(t)),$
where $Z^{\mu}:=\{x^{\mu}\}\times\mathcal{Y}^{\mu}$.
Note that Theorem 4.1 holds for any fixed $\mu>0$. Hence, combining Theorems 3.1 and 4.1, we can conclude that if $\mu$ is sufficiently small (which does not need to depend on the number of iterations $t$ or $T$), player $x$’s strategy $x^{t}$ updated by AsymP-GDA converges to a minimax strategy in the original game (1):
@@
Assume that the perturbation strength $\mu$ is set such that $\mu\in(0,\frac{\alpha}{\max_{x\in\mathcal{X}}\left\|x\right\|})$, and the learning rate is set so that $\eta\leq\frac{\mu}{\mu^{2}+\|A\|^{2}}$. Then, AsymP-GDA ensures the convergence of $x^{t}$ to an equilibrium $x^{\ast}$ in the original game at a rate of $\exp(-\mathcal{O}(t))$:
$\left\|x^{\ast}-x^{t}\right\|^{2}\leq\left\|\Pi_{\mathcal{Z^{\mu}}}(z^{1})-z^{1}\right\|^{2}\cdot\exp(-\Theta(t)).$
Corollary 4.2 provides a linear last-iterate convergence rate for AsymP-GDA. This rate is competitive with those of optimistic methods, such as OGDA and OMWU in certain settings (e.g., bilinear games) (Wei et al., 2021). Empirically, however, AsymP-GDA exhibits faster convergence in our experiments (see Figure 6 in Appendix A).
4.2 Parameter-free AsymP-GDA
Although Corollary 4.2 guarantees convergence to an equilibrium of the original game only when the perturbation strength $\mu$ is sufficiently small, the appropriate scale of $\mu$ may be unknown a priori. To eliminate this tuning requirement, we also introduce a parameter-free variant (Algorithm 1 in Appendix D) that solves asymmetrically perturbed games along a sequence of perturbation strengths $\mu\in(0,\mu_{\mathrm{init}}]$, starting from an arbitrarily large $\mu_{\mathrm{init}}>0$. For each perturbed game, the algorithm runs AsymP-GDA until the duality gap falls below a prescribed threshold, and then checks whether the resulting strategy profile achieves $\mathrm{NashConv}\leq\varepsilon$. If not, it halves $\mu$ and moves on to the next perturbed game.
Thanks to Theorem 3.1, it suffices to solve only finitely many perturbed games, regardless of any target accuracy $\varepsilon$. Since each perturbed game can be solved at a linear rate by AsymP-GDA, the overall procedure retains a linear rate to an equilibrium of the original game:
Theorem 4.3.
For any target accuracy $\varepsilon>0$ and any initialization $\mu_{\mathrm{init}}>0$, Algorithm 1 returns a strategy profile $(x,y)$ with $\mathrm{NashConv}(x,y)\leq\varepsilon$ after at most $\mathcal{O}(\ln(1/\varepsilon))$ iterations in total.
Whereas Liu et al. (2023) adopt a similar decreasing-$\mu$ mechanism, their analysis yields a $\tilde{\mathcal{O}}(1/t)$ rate. In contrast, our analysis in the above theorem shows a faster linear rate. This gap stems from the fact that our decreasing-$\mu$ procedure requires solving only a bounded number of perturbed games, while in Liu et al. (2023) the number of perturbed games grows as one targets higher accuracy.
Figure 4: Performance in extensive-form games. AsymP-DGDA performs two strategy updates per iteration for each player. For a fair comparison across methods with different per-iteration computational costs, we report the total number of strategy updates on the x-axis rather than iterations.
@@ 4.2 Parameter-free AsymP-GDA
Firstly, leveraging the strong convexity of the perturbation payoff function, $\frac{\mu}{2}\left\|x\right\|^{2}$, we can show that the distance between the current strategy profile $z^{t}=(x^{t},y^{t})$ and any equilibrium $z^{\mu}=(x^{\mu},y^{\mu})$ monotonically decreases under $\eta\leq\frac{\mu}{\mu^{2}+\|A\|^{2}}$. Specifically, we have for any $t\geq 1$:
$\left\|z^{\mu}-z^{t+1}\right\|^{2}-\left\|z^{\mu}-z^{t}\right\|^{2}$
$\!\leq\!-\!\left(\eta\mu\!-\!\eta^{2}\!\left(\mu^{2}\!+\!\|A\|^{2}\right)\right)\!\left\|x^{\mu}-x^{t+1}\right\|^{2}\!-\frac{1}{2}\left\|z^{t+1}\!-z^{t}\right\|^{2}\!.$ (7)
@@
The primary technical challenge is deriving the term related to the distance between $y^{t+1}$ and the maximin strategies set $\mathcal{Y}^{\mu}$, i.e., $\left\|\Pi_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}$, which leads to the last-iterate convergence rate. To this end, we derive the following lower bound on the path length $\left\|z^{t+1}-z^{t}\right\|$ by the distance $\|x^{\mu}-x^{t}\|^{2}$ and $\|y^{\mu}-y^{t}\|^{2}$ (as shown in Lemma E.1).
Let us represent the update rule of AsymP-GDA in (6) as $z^{t+1}=T_{\mu}(z^{t})$ using the operator $T_{\mu}:\mathcal{Z}\to\mathcal{Z}$. We show that $T_{\mu}$ is a piecewise affine mapping on $\mathcal{Z}$ (in Lemma F.3). Hence, there exists finitely many polyhedral sets such that $T_{\mu}$ can be represented as an affine function within each polyhedral set. For each set and the corresponding affine function, we can apply Hoffman’s bound (Hoffman, 1952), resulting in the following metric subregularity:
$\forall z\in\mathcal{Z},~\left\|T_{\mu}(z)-z\right\|\geq\Omega\left(\left\|z-\Pi_{Z^{\mu}}(z)\right\|\right),$
where $\mathcal{Z}^{\mu}:=\{x^{\mu}\}\times\mathcal{Y}^{\mu}$. By setting $z=z^{t}$ in the above inequality, we obtain the following lower bound on $\left\|z^{t+1}-z^{t}\right\|$ by the distance between the current strategy profile and the equilibrium set:
$\left\|z^{t+1}-z^{t}\right\|\geq\Omega\left(\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|\right).$ (8)
@@
$\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t+1})-z^{t+1}\right\|^{2}$
$\leq\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}-\Omega\left(\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}\right),$
Therefore, by mathematical induction, we finally obtain the following upper bound on the distance $\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}$:
$\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}\leq\exp\left(-\Theta(t)\right).$
@@
The main technical challenge in proving Theorem 4.1 arises from the asymmetric nature of the perturbation, which is applied only to player $x$. Unlike the symmetric case, where strong convexity in both players’ payoff functions directly yields contraction, the asymmetric setting requires a more subtle analysis since player $y$’s payoff function remains linear. Our proof establishes a linear last-iterate convergence rate by lower-bounding the variation of $z^{t}$ by the distance from $z^{t}$ to the equilibrium set $\mathcal{Z}^{\mu}$, as shown in (8), rather than relying on strong convexity. A further complication is the presence of projections onto $\mathcal{X}$ and $\mathcal{Y}$: the update mapping $T_{\mu}$ is piecewise affine, meaning that the affine mapping governing the update can switch across iterations. As a result, the dynamics cannot be analyzed as a single linear system, in contrast to unconstrained linear-quadratic games (Zhang et al., 2022). See Appendix F for details.
5 Dilated AsymP-GDA for Extensive-Form Games
We now extend our approach to extensive-form games, which model sequential decision-making under imperfect information. To cast these games as saddle-point problems, we adopt the sequence-form representation (von Stengel, 1996). In the sequence-form representation, each player’s strategy is parameterized by realization (reach) probabilities over feasible action sequences, rather than by enumerating pure strategies (i.e., complete behavior plans). Moreover, to reduce the per-iteration computational cost of AsymP-GDA in the sequence-form representation, we use the dilated Euclidean regularizer (Hoda et al., 2010) in the following experiments.
Sequence-form representation.
We consider two-player zero-sum extensive-form games with perfect recall. Let $\mathcal{I}_{x}$ and $\mathcal{I}_{y}$ denote the sets of information sets of players $x$ and $y$, respectively. For each information set $i\in\mathcal{I}_{x}$, let $A(i)$ be the set of actions available at $i$.
We use the sequence-form representation to express each player’s strategy. For player $x$, we write:
$x=(x_{i,a})_{i\in\mathcal{I}_{x},a\in A(i)},$
where $x_{i,a}\geq 0$ denotes the reach probability of taking action $a$ at information set $i$, counting only the contribution of player $x$’s own action probabilities. To express the constraints for the strategy space $\mathcal{X}$, for each information set $i\in\mathcal{I}_{x}$, we define $\mathrm{parent}(i)$ to be the unique parent (predecessor) pair [^4] The uniqueness of the predecessor is guaranteed by the perfect recall assumption. $(j,b)$ (an information set $j$ and an action $b$) on player $x$’s own decision sequence that immediately precedes $i$. We also write $\mathrm{parent}(i)=\varnothing$ if no such predecessor exists, namely if $i$ is a root information set for player $x$ along its own decision sequence. The strategy space $\mathcal{X}$ is then given by the following constraints:
$\forall i\in\mathcal{I}_{x},~\sum_{a\in A(i)}x_{i,a}=x_{\mathrm{parent}(i)}\mbox{ and }x_{i,a}\geq 0,$
where $x_{\mathrm{parent}(i)}$ denotes $x_{j,b}$ if there exists $(j,b)$ such that $(j,b)=\mathrm{parent}(i)$, and denotes $1$ if $\mathrm{parent}(i)=\varnothing$. We define $\mathcal{Y}$ analogously for player $y$.
In the sequence-form representation, the two-player zero-sum extensive-form game admits a bilinear saddle-point formulation: there exists a matrix $A$ such that $\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}x^{\top}Ay$.
Dilated AsymP-GDA.
Due to the computational efficiency of the strategy update, we propose a variant of AsymP-GDA that employs the dilated squared Euclidean regularizer (instead of the standard squared Euclidean regularizer) for both the proximal regularizer and the perturbation term. We refer to this variant as AsymP-DGDA. For comparison, we also evaluate a symmetrically perturbed counterpart, SymP-DGDA.
For player $x$, we define the dilated squared Euclidean regularizer (Hoda et al., 2010):
$\psi_{\mathrm{dil}}(x)=\frac{1}{2}\sum_{i\in\mathcal{I}_{x}}\alpha_{i}\sum_{a\in A(i)}\frac{x_{i,a}^{2}}{x_{\mathrm{parent}(i)}},$
where $\alpha_{i}>0$ is a weight parameter, and similarly for player $y$. Using $\psi_{\mathrm{dil}}$, AsymP-DGDA updates each player’s strategy as:
$x^{t+1}$ $=\mathop{\rm arg~min}\limits_{x\in\mathcal{X}}\left\{\eta\langle x,Ay^{t}+\mu\nabla\psi_{\mathrm{dil}}(x^{t})\rangle+D_{\psi_{\mathrm{dil}}}(x,x^{t})\right\},$
$y^{t+1}$ $=\mathop{\rm arg~min}\limits_{y\in\mathcal{Y}}\left\{\eta\langle y,-A^{\top}x^{t+1}\rangle+D_{\psi_{\mathrm{dil}}}(y,y^{t})\right\},$
where $D_{\psi_{\mathrm{dil}}}(a,b)=\psi_{\mathrm{dil}}(a)-\psi_{\mathrm{dil}}(b)-\langle\nabla\psi_{\mathrm{dil}}(b),a-b\rangle$. In contrast, SymP-DGDA applies perturbation symmetrically to both players. As with AsymP-GDA, AsymP-DGDA adds only negligible per-iteration runtime or memory overhead relative to standard Dilated GDA.
Empirical performance.
@@ 5 Dilated AsymP-GDA for Extensive-Form Games
We compare the NashConv of the last-iterate strategy for AsymP-DGDA against SymP-DGDA and baseline algorithms, including Dilated MWU (DMWU), Dilated GDA (DGDA) [^5] DMWU corresponds to mirror descent on the sequence-form strategy space with the dilated entropy regularizer $\psi_{\mathrm{dil\mbox{-}ent}}(x)=\sum_{i\in\mathcal{I}_{x}}\alpha_{i}\sum_{a\in A(i)}x_{i,a}\ln\frac{x_{i,a}}{x_{\mathrm{parent}(i)}}$. DGDA corresponds to mirror descent with the dilated Euclidean regularizer., Dilated OMWU (Kroer et al., 2020; Lee et al., 2021), and Dilated OGDA (Farina et al., 2019; Lee et al., 2021). Our experiments focus on five different extensive-form games: Kuhn Poker, Leduc Poker, Liar’s Dice (with four-sided), and Goofspiel (with four-card and five-card variants), all of which are implemented using LiteEFG (Liu et al., 2024). The detailed hyperparameters of the algorithms, tuned for best performance, are shown in Table 4 in Appendix A.
@@
Figure 4 shows the NashConv values for each game. As indicated by these results, AsymP-DGDA not only converges with a competitive or faster speed than any other method in all games but also directly reaches an equilibrium strategy, whereas SymP-DGDA converges near the equilibrium. These results confirm that the asymmetric perturbation leads to convergence in extensive-form games.